### v1

# Case2Code: Learning Inductive Reasoning with Synthetic Data

Abstract
Complex reasoning is an impressive ability shown by large language models (LLMs). Most LLMs are skilled in deductive reasoning, such as chain-of-thought prompting or iterative tool-using to solve challenging tasks step-by-step. In this paper, we hope to focus on evaluating and teaching LLMs to conduct inductive reasoning, that is, LLMs are supposed to infer underlying rules by observing examples or sequential transformations. However, collecting large-scale and diverse human-generated inductive data is challenging. We focus on data synthesis in the code domain and propose a Case2Code task by exploiting the expressiveness and correctness of programs. Specifically, we collect a diverse set of executable programs, synthesize input-output transformations for each program, and force LLMs to infer the underlying code implementations based on the synthetic I/O cases. We first evaluate representative LLMs on the synthesized Case2Code task and demonstrate that the Case-to-code induction is challenging for LLMs. Then, we synthesize large-scale Case2Code training samples to train LLMs to perform inductive reasoning. Experimental results show that such induction training benefits not only in distribution Case2Code performance but also enhances various coding abilities of trained LLMs, demonstrating the great potential of learning inductive reasoning via synthetic data.111Code and datasets will be available at https://github.com/choosewhatulike/case2code.
Case2Code: Learning Inductive Reasoning with Synthetic Data
Yunfan Shao1,2, Linyang Li2†, Yichuan Ma1,2, Peiji Li1,2, Demin Song2, Qinyuan Cheng1,2, Shimin Li1, Xiaonan Li1, Pengyu Wang1, Qipeng Guo2, Hang Yan2,3, Xipeng Qiu1†, Xuanjing Huang1, Dahua Lin2,3 1School of Computer Science, Fudan University 2Shanghai AI Laboratory 3The Chinese University of Hong Kong {yfshao19, xpqiu}@fudan.edu.cn {lilinyang, yanhang}@pjlab.org.cn
1 Introduction
The success of large language models (LLMs), exemplified by GPT-4 OpenAI (2023) has revolutionized the AI community. One of the most impressive abilities of LLMs is the deductive reasoning ability via chain-of-thoughts Wei et al. (2022b), exemplified by solving mathematical reasoning tasks such as GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). The major contribution to the high-level performances of reasoning problem solving is to generate and train on deductive reasoning paths via chain-of-thoughts (CoTs) using LLMs Yu et al. (2023); Mitra et al. (2024). Through various searching strategies and prompting algorithms Chen et al. (2024); Wang et al. (2023a); Yao et al. (2023); Wang et al. (2022b), LLMs can synthesize high-quality CoTs and perform deduction reasoning in many domains.
Despite the success of deducing chain-of-thought reasoning, LLMs are rarely trained to perform inductive reasoning. Inductive reasoning is a fundamental cognitive process for humans, playing a crucial role in learning, problem-solving, and scientific discovery. For instance, famous scientists make inductive judgments from actual physical phenomena, such as Newton’s Law of Motion, and Kepler’s Law based on data from Tycho. It is essential to teach artificial intelligence systems utilizing inductive reasoning to find underlying rules from facts, transformations, and logs. In this paper, to equip LLMs with such inductive reasoning ability, we introduce Case2Code, a diverse and challenging inductive reasoning task for LLMs.
Specific and limited inductive reasoning has been studied in the machine learning field. Works such as DEER Yang et al. (2022) usually design a common knowledge induction process and challenge neural models to reason inductively to find the hidden fact. Works such as DreamCoder Ellis et al. (2021) and DeepCoder Balog et al. (2016), usually synthesize a small-scale toy task to train a domain-specific model, such as learning to summarize the list operations from given list change logs.
Different from these inductive reasoning tasks, we focus on large-scale data synthesis with programs in the real world. In Case2Code, inductive reasoning samples are synthesized from real-world productive functions, which are closer to the actual distribution of general LLM applications and production. Specifically, the Case2Code challenge requires the LLM to infer the underlying program based on several input-to-output cases generated by the real-world program. In Case2Code learning, LLMs are supposed to write solutions formulated by codes based on the example outputs, which is one common scenario in the real-world working process, using examples to convey knowledge.
To obtain large-scale and diverse Case2Code data, we first gather a diverse collection of executable code texts that cover a wide range of real-world applications. Then, we generate the input-output transformation cases with the assistance of LLMs and code interpreters, which do not require powerful LLMs with advanced reasoning capabilities, resulting in a strong-to-weak distilling process. By incorporating LLMs to write input examples for each program and execute the program with these inputs to gather the corresponding outputs, we can synthesize large-scale Case2Code samples with diverse data transformations and complicated control logic.
Based on the synthetic data, we can form a unique and challenging task to evaluate and further train the LLMs and study the induction reasoning ability of LLMs. In the Case2Code challenge, we first test how current LLMs perform in making inductive reasoning. We then train LLMs with Case2Code data to further study whether such data can improve the induction reasoning ability and generalize to other commonly used reasoning tasks. Experimental results show that Case2Code is a challenging task for LLMs, even for powerful LLMs like LLaMA3-70B, GPT-3.5, and GPT-4. With constructed Case2Code data, we can boost LLMs to learn to make such inductive reasoning, while such ability can be transferred to help improve general reasoning tasks such as HumanEval and MBPP in code generation.
To summarize, in this paper, we:
(1) We introduce an induction reasoning task for LLMs, Case2Code, pointing out the necessity to synthesize inductive reasoning data for LLMs.
(2) We can teach LLMs to make induction reasoning, improving open-source LLMs by a great margin in the Case2Code challenge.
(3) Equipped with our proposed Case2Code ability, open-source LLMs can be further improved in general reasoning tasks.
2 Related Work
Our work discusses the reasoning ability of LLMs, touching on the following grounds:
2.1 Inductive Reasoning
While Reasoning is one major topic for neural networks, especially in the era of LLMs Huang and Chang (2022), inductive reasoning is rarely discussed in LLM reasoning, most research focuses on specific scenarios with limited inductive reasoning. One pioneer work is prerequisite toy tasks Weston et al. (2015) where the task goal is to solve simple induction. Later, Yang et al. (2022) introduces various world-wide knowledge such as botany, history and geography into the facts given and asks neural models to predict whether a given rule is correct. On the other hand, several works focus on training inductive program synthesis models for constrained scenarios with limited search spaces, such as operations on list, string, and manually-defined objects Balog et al. (2016); Devlin et al. (2017); Ellis et al. (2021); Shi et al. (2023). Different from previous reasoning ability studies, our proposed Case2Code task leverages diverse code in the real world as a powerful platform for LLMs to learn inductive reasoning under various challenging scenarios.
2.2 Synthetic Reasoning Data
The most widely studied LLMs reasoning task is the deductive reasoning ability, represented by chain-of-thoughts (CoTs) reasoning Wei et al. (2022b), which instructs LLMs to solve problems in detailed deductive steps. Recent works focus on building high-quality CoTs through strong LLMs such as GPT-4 to enhance smaller LLMs Yu et al. (2023); Mitra et al. (2024); Luo et al. (2023). While a particular line of work focuses on studying different search strategies of reasoning paths, including self-consistency Wang et al. (2022a), rejection sampling Huang et al. (2023); Yuan et al. (2023); Wang et al. (2023a), tree-structure CoT (ToT) searching Yao et al. (2023), Monte Carlo Tree Searching Silver et al. (2016); Chen et al. (2024), etc.
3 Method
In this section, we illustrate the framework for synthesizing Case2Code data in detail, which focuses on producing large-scale and high-quality inductive reasoning data in the code domain. Unlike other synthetic data frameworks that distill high-quality training data from a strong teacher LLM to provide supervision signals to improve student LLMs, our Case2Code synthetic framework introduces a writer LLM to assist the synthesis of data samples. Thus the overall data quality does not directly rely on the performance of the LLM generator. And we can efficiently obtain reliable Case2Code training data at scale.
3.1 Problem Formulation
The inductive reasoning task aims to find a general hypothesis based on a small set of observations to explain a phenomenon. In this paper, we define Case2Code, an inductive reasoning task in the code domain. Case2Code is a program synthesis task that targets the reconstruction of unknown programs based on observations of the program behaviors.
Formally, for a functional program , we have a set of input-output examples , where . The goal of Case2Code is to implement a program that captures the functionality of the program based on the observed set of input-output example cases . And for any new input case , the implemented program should satisfy that .
3.2 Framework Overview
In our synthetic data generation framework, we focus on generating large-scale and diverse Case2Code data automatically. As shown in Figure 2, we first collect diverse programs from large-scale datasets with rule-based filters. Then we incorporate LLMs to write diverse example inputs and utilize the code interpreter to calculate their corresponding outputs for each program. Finally, we filter out low-quality programs based on their outputs and convert the obtained triple (program, inputs, outputs) into Case2Code data for inductive reasoning in the code domain.
Note that the correctness of our synthetic data does not depend on the capabilities of the used LLMs. Therefore, we can synthetic high-quality Case2Code data at scale using small LLMs with low costs.
3.3 Collecting Programs
To obtain massive data samples for inductive reasoning learning, we first need to acquire massive and diverse programs that take input arguments, do some complicated processes, and return output values. Instead of prompting LLMs to generate functions that meet these requirements, we collect human-written high-quality programs in the wild to enhance diversity.
Specifically, we sample valid Python functions from The Stack Kocetkov et al. (2022) to construct our reasoning dataset. We incorporate the out-of-box Abstract Syntax Tree (AST) parsing tool 222https://docs.python.org/3/library/ast.html to parse each file in The Stack to obtain Python functions. We only keep self-contained high-quality functions that satisfy all of these filtering rules: (1) pass the syntax check; (2) have one or more input arguments and return values; and (3) do not rely on third-party packages or external I/O operations. After collecting these functions, we can easily execute and verify these functions to obtain diverse Case2Code data with a simple and fast code interpreter at scale, which avoids extra file or network operations that require a sophisticated sandbox.
3.4 Generating Inputs
Once we collect large-scale functions, the next step is to obtain the corresponding input-output pairs for each function to construct the Case2Code data. It is infeasible to write test cases for each function manually. So, we utilize LLMs to generate suitable input examples for these functions. We prompt LLMs to write some example input arguments for each function based on the corresponding function implementation. Detailed prompt is listed in Table 6 in the appendix.
To generate suitable input arguments, the LLM needs first to analyze the implementation of the functions, then infer the possible types and value ranges of the input arguments, and finally come up with correct input arguments. However, we argue that a powerful LLM is not the key factor for our synthetic data. As we find that while strong LLMs can write high-quality inputs to generate Case2Code training data that boosts the reasoning performance of weak LLMs, the weak LLM can also write inputs for creating Case2Code data to self-improve their reasoning ability (see Sec 4.4). Therefore, the generation process can be scaled efficiently at a low cost by using small LLMs.
3.5 Obtain Outputs
After collecting self-contained functions and the corresponding inputs, it is intuitive to incorporate a code interpreter to run these functions on their inputs for output curation. Since the LLM-generated input examples can contain errors, we introduced a filtering procedure to reject invalid inputs or functions based on their returned outputs. Specifically, if the outputs of a function do not change as the inputs change (e.g. always return the same output or exceptions), the function is considered invalid and will be filtered out.
Moreover, we also filter out functions that generate very long output values to ensure the length of the generated Case2Code data is within the context window size of current LLMs. Note that we do not filter out inputs that lead to exceptions or runtime errors, as we believe that failure call attempts can also provide valuable information for inductive reasoning to reconstruct the function.
3.6 Post-processing
The final step is to convert the obtained functions and their corresponding input-output pairs into Case2Code style data. Formally, for a given function and its test cases , we randomly sample examples () as the observed set . We generate the prompted data that facilitate the LLM to conduct inductive reasoning on the observed examples to reconstruct the given function . Converted training examples are shown in Table 7 in the appendix.
We find that the diversity of the prompts can substantially affect the generalization of the model reasoning performance (as shown in Sec 4.4). Therefore, we manually construct about 10 prompts with different styles to enhance the data diversity.
4 Experiment
In this section, we illustrate the experimental setups and discuss the experimental results to demonstrate the challenge of solving Case2Code problems and show the effectiveness of large-scale Case2Code synthetic data.
| Size | HumanEval | HumanEval+ | MBPP | MBPP+ | Case2Code | |
| GPT-4 | - | 90.2 | 86.6 | 85.7 | 73.3 | 43.6 |
| GPT-3.5 | - | 76.8 | 70.7 | 82.5 | 69.7 | 34.2 |
| LLaMA2-Chat | 7B | 14.0 | 11.6 | 26.8 | 20.3 | 0.2 |
| 13B | 23.1 | 19.5 | 37.0 | 27.6 | 8.2 | |
| 34B | 22.6 | - | 33.0 | - | - | |
| 70B | 36.6 | 28.7 | 46.3 | 35.1 | 7.8 | |
| CodeLLaMA-Instruct | 7B | 37.8 | 35.4 | 59.5 | 46.8 | 14.2 |
| 13B | 42.7 | 38.4 | 63.5 | 52.6 | 19.0 | |
| 34B | 51.8 | 43.9 | 69.3 | 56.3 | 22.6 | |
| LLaMA3-Instruct | 8B | 61.6 | 56.7 | 70.1 | 59.3 | 10.4 |
| 70B | 77.4 | 72.0 | 82.3 | 69 | 22.6 |
4.1 Experimental Setup
Data Construction
We randomly sampled about 2.3 million functions from The Stack pre-training dataset, in which we already performed data deduplication with the evaluation benchmarks (e.g. HumanEval, MBPP, etc). We conduct the data synthetic pipeline incorporating InternLM2-7b Cai et al. (2024) to generate input examples for each function. The temperature is set to 0.2 and the top_p is set to 0.95. The generation takes about 500 GPU hours using A800 GPUs. Then we use 64 CPUs to execute and filter functions, which takes about 1 hour. The execution is under a constrained Python environment to ensure safety. We eventually obtained 1.3M high-quality functions with input-output pairs for Case2Code reasoning. We hold out 500 samples for evaluation and the rest for training. For the hold-out evaluation samples, we further prompted GPT-4 (gpt-4-turbo-2024-04-09) to generate additional input examples and collect the corresponding outputs for a more strict inductive reasoning evaluation.
Training Setup
To demonstrate the generalization and effectiveness of our synthetic training data, we conduct three variants of Case2Code training: direct fine-tuning, mixed pre-training, and mixed fine-tuning. All Case2Code variants are trained for 5k steps with a batch size of 64, a maximum context window size of 4096, and apply linear warmup and cosine decay of the learning rate from the peak value of 2e-5 to 5e-6. All model training is completed on two servers of eight A800 GPUs. We conduct training on open-sourced models, i.e. InternLM2-7B Cai et al. (2024) and LLaMA3-8B AI@Meta (2024) to verify the effectiveness of synthetic training data on different model series.
Evaluation Setup
We evaluate the coding ability of trained LLMs with HumanEval, MBPP. To conduct strict evaluation, we use EvalPlus, an extension to the original HumanEval and MBPP with massive additional test cases. For models that are not instructed tuned, we apply zero-shot prompting and four-shot prompting for HumanEval and MBPP evaluation, respectively. And for instructed-aligned LLMs, we use zero-shot prompting on all these benchmarks. To evaluate inductive reasoning on code, we test various LLMs on solving Case2Code tasks, with zero-shot prompting. When evaluating the instructed models that are not tuned on Case2Code task, we find the performance is unstable and sensitive to the prompts. We manually optimized the prompts for Case2Code evaluation to elicit the actual inductive reasoning ability of these models. We use greedy decoding during the inference for all experiments.
Models
We compare the trained models with several families of representative LLMs: GPT seriesOpenAI (2023), CodeLLaMA Rozière et al. (2023), LLaMA2 Touvron et al. (2023) and LLaMA3 AI@Meta (2024). For GPT series, we evaluate GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4-turbo-2024-04-09). For other model series, we evaluate their available open-sourced versions.
Train w/
Ours |
HumanEval | HumanEval+ | MBPP | MBPP+ | Case2Code | |
| InternLM2-7B-Base | ✗ | 31.1 | 21.3 | 51.4 | 40.3 | 27.2† |
| w/ Direct Fine-tuning | ✓ | 44.5 | 34.8 | 56.0 | 40.4 | 44.4 |
| w/ Mixed Pre-training | ✓ | 43.9 | 40.9 | 58.4 | 42.6 | 41.4 |
| InternLM2-7B | ✗ | 39.0 | 33.4 | 56.8 | 54.1 | 25.6† |
| w/ Direct Fine-tuning | ✓ | 43.3 | 40.9 | 54.5 | 40.6 | 44.5 |
| w/ Mixed Pre-training | ✓ | 47.6 | 37.2 | 58.4 | 45.6 | 42.4 |
| w/ Insturction-tuning | ✗ | 49.4 | 43.9 | 58.0 | 50.4 | 6.2 |
| w/ Mixed Instruction-tuning | ✓ | 64.6 | 56.7 | 63.4 | 52.4 | 44.0 |
| LLaMA3-8B | ✗ | 35.4 | 20.1 | 59.1 | 45.1 | 29.2† |
| w/ Direct Fine-tuning | ✓ | 43.2 | 39.0 | 50.6 | 35.1 | 44.8 |
| w/ Mixed Pre-training | ✓ | 47.6 | 40.9 | 55.6 | 41.1 | 42.6 |
| w/ Insturction-tuning | ✗ | 49.8 | 45.7 | 57.6 | 47.9 | 8.6 |
| w/ Mixed Instruction-tuning | ✓ | 64.8 | 57.9 | 71.2 | 53.1 | 45.0 |
4.2 Zero-shot Case2Code Performance
As shown in Table 1, we report the zero-shot Case2Code performance of different representative LLMs and their programming performance. We can find that the zero-shot Case2Code performance of representative models is strongly related to their corresponding program synthesis performance. Models with higher program synthesis scores tend to achieve higher Case2Code performance. And larger models often outperform small models. This indicates that Case2Code can become a good benchmark to reflect the code reasoning performance of different LLMs. However, the zero-shot Case2Code scores of LLMs have a large gap compared with their coding accuracy, which demonstrates that existing LLMs are better at some types of reasoning (e.g. writing programs based on instructions) than others (e.g. inductive programs by their behaviors). This can be explained as the LLMs are trained with massive program generation data but fewer samples similar to Case2Code that need inductive reasoning. Similar to the Reverse Curse Berglund et al. (2023), models trained with deductive reasoning data struggle to transfer to inductive reasoning tasks.
4.3 Generalization of Case2Code
One essential issue of synthetic data is its generalization ability. Therefore, we train different LLMs with our synthetic Case2Code dataset under various settings to explore how it affects the learning of code reasoning of LLMs.
4.3.1 Direct Fine-tuning
First, we find that LLMs that are directly trained on the Case2Code reasoning samples can effectively learn coding based on cases. As shown in Table 2, by direct fine-tuning, Internlm2-7B and LLaMA3-8B can significantly outperform the few-shot prompting baselines by up to 18.9%, achieve up to 44.5% and 42.0% accuracy on Case2Code evaluation set, respectively, which even outperforms the more powerful LLMs like LLaMA3-70B, GPT-3.5, and comparable with GPT-4 (results in Table 1). Moreover, models trained with Case2Code reasoning also improve their program synthesis performance on benchmarks like HumanEval and MBPP. This indicates that the Case2Code reasoning is general and challenging. Training on Case2Code samples not only boosts the inductive reasoning performance in distribution but enhances the code understanding and code generation abilities of LLMs. As the Case2Code samples can be synthetic at scale, we believe that synthesizing large-scale and high-quality inductive reasoning data is a promising path to consistently improve LLMs without exhausting data.
4.3.2 Mixed Training
Then, we explore how to better incorporate our synthetic Case2Code data into different stages of LLM training to enhance the reasoning ability of LLMs in general. Specifically, we train LLMs with two variants of data mixing, either during pre-training or in the supervised fine-tuning (SFT) stage. The first mixing strategy introduces natural language pre-training texts from the Pile Gao et al. (2021) and the code pre-training samples from The Stack Kocetkov et al. (2022). The mixing ratio is 1:1:2 for samples from the Pile, The Stack, and the Case2Code dataset, respectively. On the other hand, we incorporate a supervised fine-tuning (SFT) dataset from WizardCoder Luo et al. (2023) to demonstrate that the performance gain of Case2Code training does not come from the understanding of instructions but the learning of inductive reasoning of code execution. We combine the SFT dataset with Case2Code samples in a 1:3 ratio, as the size of our synthetic dataset is much larger.
Mixed Pre-training
As shown in Table 2, when incorporated into the pre-training stage, the Case2Code training data helps the model to connect the execution states with the function implementation, which further facilitates the program synthesis performance of these LLMs. Compared with directly fine-tuned on Case2Code dataset, training these samples with pre-training texts enables the generalization of inductive reasoning of code states learned by the Case2Code task.
Mixed Instruction-tuning
When trained with instruction-following datasets, the Case2Code data also improves the performance of the programming with instruction tasks, as reported in Table 2. We evaluate the SFT models with the zero-shot instructed version of programming synthesis tasks, HumanEval, and MBPP. We find that incorporating Case2Code data boosts the performance of various LLMs on code generation tasks. Compared to the corresponding SFT baselines, InternLM2-7B improves on HumanEval from 49.4% to 64.6%, with more than 10% improvements. LLaMA3-8B achieves 64.6%, 57.9%, and 71.2% on HumanEval, HumanEval+, and MBPP, respectively, with significant improvements compared to the SFT version. These results demonstrate the effectiveness of learning on Case2Code and the necessity of incorporating inductive reasoning data into LLM training.
4.4 Ablation Study
In this section, we conduct ablation studies to demonstrate the effectiveness of the Case2Code synthetic pipeline across different families and scales of LLMs.
Prompt Diversity
Since the synthetic Case2Code training data is converted by triples of (programs, inputs, outputs), during the construction, the prompt templates are utilized to embed the input-output pairs to form natural language texts for LLM to learn. As the LLM can only rely on these converted prompts to learn the Case2Code, it is important to understand the effectiveness of how different prompt templates affect the training of LLMs. Intuitively, the diversity of prompt templates plays an important role in the learning of LLMs. Therefore, we compare synthetic data prompted using a single template style with data utilizing diverse styles of templates. The result is reported in Figure 3, in which diverse prompts may have little effect on the in-domain Case2Code performance, however, the diversity significantly affects the accuracy of LLMs on out-of-domain program synthesis tasks. It is indicated that diversity can be critical during LLM learning, which also has been discussed in other domains like in general natural language processing tasks Wei et al. (2022a) and alignment Ouyang et al. (2022); Wang et al. (2023b).
| TP | TGS | Costs | # Samples | |
| InternLM2-7B | 1 | 1600 tokens/s | 1 | 1.3M |
| LLaMA3-70B | 4 | 720 tokens/s | 4.5 | 700K |
| HumanEval | HumanEval+ | MBPP | MBPP+ | Case2Code | |
| InternLM2-1.8B | 32.3 | 29.9 | 43.6 | 24.3 | 27.8 |
| InternLM2-7B | 64.6 | 56.7 | 63.4 | 52.4 | 42.2 |
| InternLM2-20B | 73.1 | 65.2 | 77.4 | 55.4 | 46.0 |
LLM for Generating Inputs
During the synthesis of Case2Code data, a critical step is prompting the LLM to write several input examples for each program. These inputs are then executed with the corresponding programs one by one to obtain the program outputs, thus we can utilize these important contexts to construct Case2Code training data. To explore whether the reasoning ability of the LLM writer affects the synthetic data quality, we replace the LLM generator from Interlm2-7B to LLaMA3-70B, and rerun the data synthesis pipeline to obtain a new version of Case2Code training data. Due to the high costs of LLaMA3-70B, we only generate half the size of our original synthetic data. Detailed generation costs are reported in Table 3. We train Interlm2-7B with this version of Case2Code dataset under the instruction-tuning setup to evaluate the data quality. As shown in Figure 4, compared with the InternLM2-7B generator, large LMs like LLaMA3-70B can write high-quality input samples that help trained LLMs to achieve comparable code reasoning capability with fewer training data. It indicates that the input generation step can affect the overall synthetic data quality, suggesting data collectors choose a strong LLM to be the input writer if compute resources are sufficient. However, we note that LLaMA3-70B contains too many parameters that are more costly than InternLM2-7B. By generating inputs with InternLM2-7B, our Case2Code data synthesis framework maintains generation efficiency and data quality. It also demonstrates the possibility of self-improving for LLMs on their code reasoning capabilities.
Model Scale
We want to explore whether the Case2Code data synthesized using a small model can still improve a large model, and how the model scale affects the learning process. Therefore, we use Case2Code data generated with InternLM2-7B to train models in the InternLM2 series to investigate these questions. The training is taken under the setting of data mixing with SFT dataset Luo et al. (2023) and the results are shown in Table 4. Our synthetic data consistently enhances the code reasoning performance of various sizes of LLMs, even though one of the student models is almost three times larger than the model used for data synthesis. These results demonstrate the possibilities of weak-to-strong supervision in code-related tasks at scale.
5 Conclusion
We first construct a new benchmark Case2Code to evaluate the inductive reasoning capability of LLMs in the code domain. Then, we propose a data synthetic framework to construct Case2Code training samples at scale. By just using small LLMs and a code interpreter, we can collect high-quality Case2Code data from pre-training code texts automatically and efficiently. By training on various LLMs in multiple settings, we demonstrate the Case2Code can improve not only the inductive reasoning ability of LLM but also the general coding capabilities. We believe synthetic Case2Code is a promising way to continue improving the LLMs when human-generated data is exhausted.
Limitations
In this work, we study Case2Code, a synthetic task for learning inductive reasoning capabilities. Our work is still limited in several aspects:
Potential harmful programs: we gather and filter programs from the pre-training code corpus, which excludes code that may contain dangerous operations like system calls, file manipulation, and network traffic that require careful safety checks and vulnerability mitigation. In the future one can incorporate a safe and reliable execution environment that supports these operations for Case2Code synthesis.
Programming languages: we focus on synthesizing Case2Code data using Python programs, as it is a commonly used programming language and can be easily and reliably manipulated and executed. Future work can extend the data synthesis framework to more programming languages and applications.
Long context: some inputs or outputs of the given programs can be extremely long, which can be challenging to fit into the context window of current LLMs. Future work can explore efficient methods of representing and learning long-context case-to-code induction.
Data modality: we represent cases in our Case2Code data as texts for LLM training, however, real-world programs often interact with multi-modal inputs and outputs like audio, image, and video. How to effectively collect and learn multi-modal inductive reasoning remains a big challenge.
Appendix A Prompts Used in Case2Code
We demonstrate the prompts used during the Case2Code synthesis, training, and evaluation as follows:
| Prompt Template for Zero-shot Case2Code Evaluation. |
| {spverbatim} prompt Please write the correct names of arguments. As the function you implement will be called by: func_name(**input_dict). Keep the original type. No need to convert the output to string. |
| Prompt for LLM Input Generator |
| {spverbatim} Given the function, first analyze the types of the function arguments, then write 10 different example inputs for the function, each example should be a dict with function arguments’ names and their values. Output format: “‘python examples = [ dict(argname=argvalue), …. ] “‘ Function: “‘python def test_func(a: int, b: str) -> str: return str(a) + b “‘ Examples: “‘python examples = [ dict(a=1, b=’a’), dict(a=2, b=’b’), dict(a=3, b=’c’), dict(a=4, b=’d’), dict(a=5, b=’e’), dict(a=6, b=’f’), dict(a=7, b=’g’), dict(a=8, b=’h’), dict(a=9, b=’i’), dict(a=10, b=’j’), ] “‘ Function: “‘python code “‘ Examples: |
| Case2Code Examples |
| {spverbatim} Input: dict(s="abcba", center=2), Output: "(5, 0, 4)" Input: dict(s="abcdefg", center=3), Output: "(1, 3, 3)" Input: dict(s="aba", center=1), Output: "(3, 0, 2)" Input: dict(s="racecar", center=3), Output: "(7, 0, 6)" Input: dict(s="madam", center=2), Output: "(5, 0, 4)" Input: dict(s="abcabcabc", center=4), Output: "(1, 4, 4)" Input: dict(s="xyzyx", center=2), Output: "(5, 0, 4)" Input: dict(s="hello", center=2), Output: "(1, 2, 2)" Input: dict(s="ab", center=0), Output: "(1, 0, 0)" Input: dict(s="a", center=0), Output: "(1, 0, 0)" Write a function that generates the output from the input. Function: greatest_palindrome_size_odd |
| {spverbatim} The function is: “‘python def greatest_palindrome_size_odd(s, center): right = center + 1 left = center - 1 size = 1 optimum_right = optimum_left = center while left >= 0 and right < len(s): if s[left] == s[right]: size += 2 optimum_left = left optimum_right = right right += 1 left -= 1 else: break return size, optimum_left, optimum_right “‘ |
| {spverbatim} Arguments and results: Input: dict(seq="ATCG", complementarity=’A’: ’U’, ’T’: ’A’, ’C’: ’G’, ’G’: ’C’), Output: "’CGAU’" Input: "ATCG", ’A’: ’T’, ’T’: ’A’, ’C’: ’G’, ’G’: ’C’, Output: "’CGAT’" Input: seq:"ACGT", complementarity:’A’: ’U’, ’T’: ’A’, ’C’: ’G’, ’G’: ’C’, Output: "’ACGU’" Input: "ACGT", ’A’: ’T’, ’T’: ’A’, ’C’: ’G’, ’G’: ’C’, Output: "’ACGT’" Please write a function to process the input arguments and produce the specified outputs. Start with the function: reverse_complement The function is: “‘python def reverse_complement(seq, complementarity): bases = list(seq) bases = [complementarity[base] for base in bases] reversed_complement = ”.join(bases) return reversed_complement[::-1] “‘ |